Find four points contained in the inverse. Express your values as an integer or simplified fraction.

unnamed (2).jpgFind four points contained in the inverse. Express your values as an integer or simplified fraction.
Answer #1

Answer: (0, 1/2), (1, 1/4), (2, 0), (3, -1/4)

Proof:

To find the inverse function, swap x and y

x = -4y + 2

Then solve for y

y = -x/4 + 1/2

Now we can plug in any 4 inputs. 0, 1, 2, and 3 are easy

-(0)/4 + 1/2 = 1/2

-(1)/4 + 1/2 = 1/4

-(2)/4 + 1/2 = 0

-(3)/4 + 1/2 = -1/4

Then combine the x and y values into coordinate pairs

(0, 1/2), (1, 1/4), (2, 0), (3, -1/4)

answered by: Guest